OpenStudy (anonymous):

I am in need of major help with these math problems. If anyone can help me , i will be really thankful ! csc2x = 1/2Csc(x)Sec(x) sin(x)/1-cos + 1-cos/sin(x) = 2Csc Sec^2x/Cot(x) - Tan^2 = Tan These problems are all "prove the identity" Thank you to everyone who participates !

jimthompson5910 (jim_thompson5910):

I'll do the first one to get you started

jimthompson5910 (jim_thompson5910):

csc(2x) = (1/2)*csc(x)*sec(x) 1/(sin(2x)) = (1/2)*csc(x)*sec(x) 1/(2*sin(x)*cos(x)) = (1/2)*csc(x)*sec(x) (1/2)*(1/sin(x))*(1/cos(x)) = (1/2)*csc(x)*sec(x) (1/2)*csc(x)*sec(x) = (1/2)*csc(x)*sec(x) So the identity is confirmed.

jimthompson5910 (jim_thompson5910):

I used the identities found on this page http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

Thank you ! I understand the identities now. They are just somewhat frustrating because of the many combinations

jimthompson5910 (jim_thompson5910):

yeah it takes a while to get used to it, but you'll get better with more practice